find an equation of the line having the specified slope and containing the indicated point. write your final answer as a linear function in slope intercept form. Then graph the line.
The problem is m=negative one fifth (-2,1)
this is what i have done and am not sure what to do from here
m=-4 (x=-1 and y=5)
y =-1times -4 +b
y= -4+ b
5=-4+9
x and y chart of x being -4 and y bring 5

Answer by richwmiller(17219)

I have no idea what you were trying to do.
We want an equation that looks something like y=2x+4
but with different numbers
It is modeled after y=mx+b where m is the slope and b is the value of y when x=0
So we are told that m=-1/5
and we have the point (-2,0)
so we need to find b
0=-1/5*(-2)+b
0=2/5+b
-2/5=b
So our equation is y=-1x/5-2/5
We could get rid of the fractions but we are told to leave it this way.
so we know the point (0,-2/5) and the point given us (-2,0)
The slope m tells us that for every 1 that y goes up x goes down 5
It also tells us that for every 1 that y goes down x goes up 5.
So from point (-2,0) we can find a point (-7,1) or (3,-1)